# Observable Primordial Gravitational Waves from Non-minimally Coupled $R^{2}$ Palatini Modified Gravity

### v1

@@ Abstract
We probe the spectrum of primordial gravitational waves (GWs) produced during the eras of hyperkination, kination, and reheating in a non-minimally coupled, $\mathcal{L}\propto(1+\xi\chi/M_{\text{Pl}})^{t}(R+\alpha R^{2})$, modified gravity using the Palatini formulation. We consider a runaway potential, which gives an era of kinetic domination after the end of inflation. The inclusion of $R^{2}$ in Palatini gives non-canonical quadratic kinetic terms, which give the phase of hyperkination. The coupling order $t$ is varied to examine a large class of theories up to $\chi^{2}R^{2}$. For models with $t>0$, reheating is not achieved naturally; hence, we supplement such theories with a reheating mechanism based on the interaction of inflaton and radiation produced at the end of inflation due to cosmological expansion. We demonstrate that the energy density of the GWs is enhanced as a function of the coupling during kination for all considered theories, and a short-lived phase of hyperkination truncates the boost and avoids the over-production of GWs. Hyperkination, and thus the $R^{2}$ term, should be deemed necessary in all theories with a runaway potential as it prevents the GW enhancement during kination from destabilizing the Big Bang Nucleosynthesis. The spectrum remains flat for the period of hyperkination and reheating. We examine the available parameter space for which the theories remain valid and place bounds on the Hubble parameter ($H$) and radiation energy density ($\Omega_{r}^{\text{end}}$) at the end of inflation. We find that as we decrease the order of the coupling, the spectra shift towards a more observable regime of future GW experiments. The observation of the plateau during reheating will constrain the $H$ and $\Omega_{r}^{\text{end}}$ values, while the spectral shape of the boost obtained during kination will confirm the nature of the theory. The bounds from hyperkination lie in the kHz-GHz frequency range whose detection can be positively anticipated via resonant cavities. If observed, it will confirm an important prediction of inflation and the existence of a modified theory of gravity.
@@ 1 Introduction
Figure 1: Evolution of the inflaton potential as a function of a field value in conventional (Left) and alternate (Right) history models.
@@
Another motivation comes from astrophysical and cosmological observations. No known theoretical models have been able to explain the observations. The simplest model that currently fits the data is the concordance model, $\Lambda$CDM. To name a few, issues like the presence of dark energy, cosmic acceleration, the existence of dark matter, and the large-scale structure of the universe came up from observations that were not explained by GR. The search for theories that explain these observations and also meet theoretical expectations led to the rise of the domain of modified gravity. Among a variety of modifications considered [41, 42, 43, 44, 45, 46, 47, 48, 49, 50] one class of theory that strikes out is the $f(R)$ theories of gravity [51, 52]. It simply generalizes the Einstein-Hilbert action as a function of the Ricci scalar,
@@
One of the key signatures from inflation and the period after it are primordial GWs, primordial because they are old. These are a direct result of the accelerated expansion of the universe and are produced as a consequence of tensor perturbations of the spacetime metric and constitute a stochastic gravitational-wave background (SGWB). An indirect artifact of these GWs is the imprint on CMB in the form of B-mode polarizations, which can be observed [63]. While these GWs can also be observed directly through interferometers, like the LIGO (Laser Interferometer Gravitational-Wave Observatory) [64, 65], Virgo [66], and KAGRA [67], and experiments like the NANOGrav (North American Nanohertz Observatory for Gravitational Waves) [68, 69], PTAs (Pulsar Timing Arrays) [70] like the European PTA (EPTA) [71, 72], the International PTA (IPTA) [73, 74], and the Parkes PTA (PPTA) [75], they are too faint to detect with our current detection capabilities. However, with future experiments like LISA (Laser Interferometer Space Antenna) [76, 77, 78, 79], ET (Einstein Telescope) [80], CE (Cosmic Explorer) [81], BBO (Big Bang Observatory) [82, 83, 84], SKA (Square Kilometer Array) [85], DECIGO (DECi-hertz Interferometer Gravitational-wave Observatory) [86, 87, 88], ASTROD-GW (Astrodynamical Space Test of Relativity using Optical Devices optimized for Gravitational Wave) [89], $\mu$Ares [90], AION-Km (Atom Interferometer Observatory and Network) [91] and AEDGE (Atomic Experiment for Dark Matter and Gravity Exploration) [92] such observations can be positively anticipated. Detection of ultra-high frequency GWs through resonant cavities also shows promise [93]. If observed, this will allow us to probe the early universe at very high energies and beyond the last scattering surface.
@@
In this paper, we try to probe the spectral energy density today of the primordial gravitational waves generated during the epochs of hyperkination, kination, and reheating. Working in the Palatini formalism, we examine a modified gravity of the form $\mathcal{L}\propto(1+\xi\chi/M_{\text{Pl}})^{t}(R+\alpha R^{2})$, where $\chi$ is our inflaton field, $R$ is the Ricci scalar, and $t$ is the variable order of our coupling. The order parameter $t$ let us probe lower-order theories ($t<0$) [94], the Starobinsky model ($t=0$) [7] up to $\chi^{2}R^{2}$ ($t=2$). We work with a non-oscillatory, runaway inflaton potential (see Figure 1) that allows the field to roll down freely after inflation. Starting with the modified gravity action in the Jordan frame, we move to the Einstein frame through a Weyl transformation to obtain non-canonical quartic kinetic terms. The quartic kinetic terms are a unique signature of $R^{2}$ and are responsible for the period of hyperkination, which can occur just after inflation and before kination. We then solve for the dynamics of the field during the epochs of kination and hyperkination. To counter the issue of reheating in this model with non-oscillatory potential, we supplement a reheating mechanism. The radiation produced at the end of inflation modifies the Lagrangian by coupling to the inflaton field, causing decay of scalar field energy density and bringing out radiation domination and, thus, a successful reheating. We calculate the tensor perturbations and their spectrum analytically during all phases of hyperkination, kination, reheating, and beyond. We examine the available parameter space for which the theories remain valid and place bounds on the Hubble parameter ($H$) and radiation energy density ($\Omega_{r}^{\text{end}}$) at the end of inflation. We explore the observability of these spectra against the sensitivity of the current and future GW detection experiments. A successful detection, if realized in the future, will strongly hint towards the presence of higher-order curvature terms and a modified theory of gravity.
@@
The paper is organized as follows: In Section 2, we discuss the mathematical formulation of the model with an overview of the action, equation of motion, and the periods of inflation, kination, and hyperkination. In Section 3, we propose a reheating mechanism that supplements our model. We provide an insight into background solutions and gravitational wave mode functions and compute their spectrum in Section 4, while Section 5 contains our results on the observability of primordial gravitational waves by current and future experiments. We conclude the paper with a brief summary in Section 6. We provide details of our calculations and derivations in Appendix A - Appendix D.
@@ 2.1 The Action and the Potential
where $\tilde{f}$ represents derivative with respect to the auxiliary field $\phi$. Note that the $\phi^{2}$ field does not have any kinetic term present, so varying the action with $\phi^{2}$ will not result in any equation of motion but rather a constraint equation.
@@ 2.2 Equation of Motion
The least-action principle demands that the variation of action with respect to the metric field is zero, $0=\delta\textbf{S}$ and since the variation holds for any $\delta g^{\mu\nu}$, it implies,
@@ 2.3 Inflation
However, we find a model, $V=\beta\chi^{2}$ and $h=\gamma\chi^{2}$, discarded by [96], as valid and claim that successful inflation is observed in $\chi^{2}R^{2}$ theory with a quadratic potential. In addition to being consistent with Planck observations [97], the model can also be consistent with sub-Planckian inflation and with potential corrections from quantum gravity whilst conserving unitarity during inflation and reheating to a sufficient temperature for a successful post-inflation cosmology [98, 57]. We can assume that inflation is successfully achieved for all orders of couplings with runaway potential as the behavior of the potential is the same for all couplings before the end of inflation.
@@ 2.4 Hyperkination and Kination
$3H^{2}=8\pi G\rho_{\chi}=\kappa\rho_{\chi}$ (2.26)
where $H$ is the Hubble parameter. Following the methodology given by [99], in order to eliminate $H$ from the equation of motion, we change the time variable to the number of $e$-folds, $N=\ln{a}$, where $\mathrm{d}N=H\mathrm{d}\tau$. We normalize the scale factor such that, for $a=e^{N}=1$ for $N=0$, $N$ denotes the number of $e$-folds after the end of inflation. Let the overhead bar denote the derivative with respect to $N$ [^2] Note that the prime on $h$ indicates the derivative with respect to $\chi$. and WLOG let us assume $\dot{\chi}>0$. From (2.26),
@@
We compare the convergence of the derivative of the field with respect to the number of $e$-folds $N$ to the kinetic attractor value for (2.42) and for different exponents of coupling $t$ from (2.32), considering all the constraints on initial conditions and length of hyperkination and transition in Figure 2. The three black dashed vertical lines denote the start of kination, $N_{\text{kin}}$, the end of inflation, and the start of hyperkination, $N_{\text{inf}}^{{\text{end}}}=0,$ and the limiting case for length of hyperkination, $N_{\text{hyp}}^{{\text{end}}}=0.2$. The constraint on transition length between hyperkination and kination from (2.48) suggests that the kinetic terms start to dominate slightly before the inflation completely ends and hyperkination lasts for a small duration. The end of hyperkination is followed by the phase of standard kination, where the field rolls forward freely. The epoch of hyperkination is fastest for the initial approximation in (2.42). The length increases as we increase the coupling exponent in (2.32), and convergence is achieved at values mentioned in Table 1. The field derivative for $t=3$ never approaches the attractor value.
@@ 3 Reheating Mechanism
where $b$ and $n$ are small and positive coupling constants. For simplicity, we assume that the bare masses of the fields $\zeta$ and $\psi$ are very small, such that $m_{\zeta}(\chi)=g\chi$ and $m_{\psi}(\chi)=|h\chi|$. We can consider the effective potential of the scalar field after kination as,
@@
Typically, coupling constants of the interaction of the inflaton field with matter are extremely small [110], we consider $b\sim\mathcal{O}(10^{-1})$ and $n\sim\mathcal{O}(10^{-2})$. $\lambda$ is the coupling constant for the Higgs-like potential. Considering it is stable, we choose $\lambda\sim\mathcal{O}(10^{-2})$. The only free parameter now is the mass of the scalar field.
@@
To get an era of reheating, the mass of the inflaton field should be fine-tuned such that the decay rate causes the inflaton field to decay faster than radiation. This happens for $m_{\chi}=10^{4}-10^{5}$ GeV, see the left panel of Figure 3. If no decay process is considered (black dashed line), the scalar field energy density will never be sub-dominant than radiation. As we increase the mass parameter, the energy density dilutes faster, eventually letting radiation take over for $m_{\chi}>8\times 10^{3}$ GeV. Considering that the decay processes take time and do not happen as quickly as for $m_{\chi}=10^{5}$ GeV, we choose the mass of the scalar field to be $10^{4}$ GeV. The length of hyperkination is fixed to the upper bound of $\Delta N_{\text{hyp}}=0.2$; for a choice of $m_{\chi}=10^{4}$ GeV, the length of the kination phase is $\Delta N_{\text{kin}}=4$. The panel on the right of Figure 3 is the new evolution of energy density accounting for the decay process in our alternate history theory for couplings with exponents $t>0$.
@@ 4.1 Background Solutions
where $\epsilon$ is the first Hubble slow-roll parameter. Note that we will denote the Hubble parameter at the end of inflation $H_{\text{end}}$ as $H$. Solving for all the epochs, see Appendix B for detailed derivation, we get
@@ 4.2 Gravitational Wave Mode Functions
where prime denoted derivative with respect to the conformal time $\eta$. From the previous section, we can write,
@@
where $H_{\nu}^{(1)}$ and $H_{\nu}^{(2)}$ are Hankel functions of the first and second kind, respectively. We solve for the mode functions in detail in Appendix C and note the final results here,
@@
To determine the constants $\alpha_{\pm}$, $\beta_{\pm}$, and $\gamma_{\pm}$; we demand the continuity of $v_{k}^{s}$ and its derivative at the transition times $\eta_{\text{end}}$, $\eta_{\text{kin}}$, $\eta_{\text{reh}}$. The details for this mode-matching process are mentioned in Appendix D. The results are as follows,
@@
We check the time-evolution of the perturbations by fixing the length of the period of hyperkination and varying the order of coupling; see Figure 5. We keep $\Delta N_{\text{hyp}}=6.18$ so that we can gauge the changes if there are any from the variation. The order of the coupling decreases as we go from top to bottom. Similar to the results obtained in 4 for kination, we see that the perturbations get a boost in energy for all orders. However, as the order decreases, the amplitude of produced gravitational wave modes during hyperkination increases, indicating that the hyperkination phase will contribute significantly to the primordial background for such couplings. Another feature is the increasing length of the period of kination with decreasing order. This suggests that the energy from GWs during kination will also increase for every decreasing order. In Figure 6, we vary $\Delta N_{\text{hyp}}$ for the $t=2$ coupling. We do not observe any change in GW perturbations generated during hyperkination. However, the kination phase reduces in length, and the modes decay faster with decreasing length of hyperkination.

### v2

@@ Abstract
We probe the spectrum of primordial gravitational waves (GWs) produced during the eras of hyperkination, kination, and reheating in a non-minimally coupled, $\mathcal{L}\propto(1+\xi\chi/M_{\text{Pl}})^{t}(R+\alpha R^{2})$, modified gravity model using the Palatini formulation under a runaway inflaton potential. The coupling order $t$ is varied to examine a large class of theories up to $\chi^{2}R^{2}$. For models with $t>0$, reheating is not achieved naturally; hence, we supplement such theories with a reheating mechanism based on the interaction of inflaton and radiation produced at the end of inflation due to cosmological expansion which gives successful radiation domination with $T_{\text{reh}}\sim 10^{8}$ GeV. We demonstrate that the energy density of the GWs is enhanced as a function of the coupling during kination for all considered theories, and a short-lived phase of hyperkination, a result of $R^{2}$ in Palatini, truncates the boost and avoids the over-production of GWs. Our work provides another strong hint towards inclusion of $R^{2}$ terms. The spectrum remains flat for the period of hyperkination and reheating. We find that as we decrease the order of the coupling, the spectra shift towards a more observable regime of future GW experiments. We also highlight the observability of GWs from minimal Starobinsky model and availability of wide parameter space. The observation of the plateau during reheating will constrain the $H$ and $\Omega_{r}^{\text{end}}$ values, while the spectral shape of the boost obtained during kination will confirm the nature of the theory. The bounds from hyperkination lie in the kHz-GHz frequency range. If observed, it will confirm an important prediction of inflation and the existence of a modified theory of gravity.
@@ 1 Introduction
Figure 1: Evolution of the inflaton potential as a function of a field value in conventional (Left) and alternate (Right) history models. $\phi_{\text{CMB}}$ marks the CMB emission, $\phi_{\text{end}}$ the end of inflation and start of hyperkination, $\phi_{\text{kin}}$ the end of hyperkination and start of kination, $\phi_{\text{rad}}$ the end of kination and start of radiation domination, $\phi_{\text{eq}}$ denotes the radiation-matter equality and the onset of matter domination while $\phi_{\Lambda}$ highlights the start of an era dominated by dark energy.
@@
Another motivation comes from astrophysical and cosmological observations. No known theoretical models have been able to explain the observations. The simplest model that currently fits the data is the concordance model, $\Lambda$CDM. To name a few, issues like the presence of dark energy, cosmic acceleration, the existence of dark matter, and the large-scale structure of the universe came up from observations that were not explained by GR. The search for theories that explain these observations and also meet theoretical expectations led to the rise of the domain of modified gravity. One such class of theory that strikes out is the $f(R)$ theories of gravity [41, 42, 43, 44]. It simply generalizes the Einstein-Hilbert action as a function of the Ricci scalar,
@@
One of the key signatures from inflation and the period after it are primordial GWs, primordial because they are old. These are a direct result of the accelerated expansion of the universe and are produced as a consequence of tensor perturbations of the spacetime metric and constitute a stochastic gravitational-wave background (SGWB). Primordial GWs were considered widely in modified gravity theories [55, 56, 57]. An indirect artifact of these GWs is the imprint on CMB in the form of B-mode polarizations, which can be observed [58]. While these GWs can also be observed directly through interferometers, like the LIGO (Laser Interferometer Gravitational-Wave Observatory) [59, 60], Virgo [61], and KAGRA [62], and experiments like the NANOGrav (North American Nanohertz Observatory for Gravitational Waves) [63, 64], PTAs (Pulsar Timing Arrays) [65] like the European PTA (EPTA) [66, 67], the International PTA (IPTA) [68, 69], and the Parkes PTA (PPTA) [70], they are too faint to detect with our current detection capabilities. However, with future experiments like LISA (Laser Interferometer Space Antenna) [71, 72, 73, 74], ET (Einstein Telescope) [75], CE (Cosmic Explorer) [76], BBO (Big Bang Observatory) [77, 78, 79], SKA (Square Kilometer Array) [80], DECIGO (DECi-hertz Interferometer Gravitational-wave Observatory) [81, 82, 83], ASTROD-GW (Astrodynamical Space Test of Relativity using Optical Devices optimized for Gravitational Wave) [84], $\mu$Ares [85], AION-Km (Atom Interferometer Observatory and Network) [86] and AEDGE (Atomic Experiment for Dark Matter and Gravity Exploration) [87] such observations can be positively anticipated. Detection of ultra-high frequency GWs through resonant cavities also shows promise [88]. If observed, this will allow us to probe the early universe at very high energies and beyond the last scattering surface.
@@
In this paper, we try to probe the spectral energy density today of the primordial gravitational waves generated during the epochs of hyperkination, kination, and reheating. Working in the Palatini formalism, we examine a modified gravity of the form $\mathcal{L}\propto(1+\xi\chi/M_{\text{Pl}})^{t}(R+\alpha R^{2})$, where $\chi$ is our inflaton field, $R$ is the Ricci scalar, and $t$ is the variable order of our coupling. $R^{2}$ models in Palatini formalism have been widely studied in the context of inflation and reheating [89, 90, 91, 92, 93]. The order parameter $t$ let us probe lower-order theories ($t<0$) [94], the Starobinsky model ($t=0$) [7] up to $\chi^{2}R^{2}$ ($t=2$). We work with a non-oscillatory, runaway inflaton potential (see Figure 1) that allows the field to roll down freely after inflation. Starting with the modified gravity action in the Jordan frame, we move to the Einstein frame through a Weyl transformation to obtain non-canonical quartic kinetic terms. The quartic kinetic terms are a unique signature of $R^{2}$ and are responsible for the period of hyperkination [95], which can occur just after inflation and before kination. We then solve for the dynamics of the field during the epochs of kination and hyperkination. To counter the issue of reheating in this model with non-oscillatory potential, we supplement a reheating mechanism. The radiation produced at the end of inflation modifies the Lagrangian by coupling to the inflaton field, causing decay of scalar field energy density and bringing out radiation domination and, thus, a successful reheating. We calculate the tensor perturbations and their spectrum analytically during all phases of hyperkination, kination, reheating, and beyond. We examine the available parameter space for which the theories remain valid and place bounds on the Hubble parameter ($H$) and radiation energy density ($\Omega_{r}^{\text{end}}$) at the end of inflation. We explore the observability of these spectra against the sensitivity of the current and future GW detection experiments. A successful detection, if realized in the future, will strongly hint towards the presence of higher-order curvature terms and a modified theory of gravity.
@@
The paper is organized as follows: In Section 2, we discuss the mathematical formulation of the model with an overview of the action, equation of motion, and the periods of inflation, kination, and hyperkination. In Section 3, we propose a reheating mechanism that supplements our model. We provide an insight into background solutions and gravitational wave mode functions and compute their spectrum in Section 4, while Section 5 contains our results on the observability of primordial gravitational waves by current and future experiments. We conclude the paper with a brief summary in Section 6. We provide details of our calculations and derivations in the appendices.
@@ 2.1 The Action and the Potential
where $\tilde{f}$ represents derivative with respect to the auxiliary field $\phi$ and $\tilde{\tilde{f}}$ represents double derivative with respect to $\phi$. Note that the $\phi^{2}$ field does not have any kinetic term present, so varying the action with $\phi^{2}$ will not result in any equation of motion but rather a constraint equation.
@@ 2.2 Equation of Motion
The least-action principle demands that the variation of action with respect to the metric field is zero, $\delta\textbf{S}=0$ and since the variation holds for any $\delta g^{\mu\nu}$, it implies,
@@ 2.3 Inflation
However, we find a model, $V=\beta\chi^{2}$ and $h=\gamma\chi^{2}$ (more in Appendix B), as valid and claim that successful inflation is observed in $\chi^{2}R^{2}$ theory with a quadratic potential. In addition to being consistent with Planck observations [98], the model can also be consistent with sub-Planckian inflation and with potential corrections from quantum gravity whilst conserving unitarity during inflation and reheating to a sufficient temperature for a successful post-inflation cosmology [99, 49]. We can assume that inflation is successfully achieved for all orders of couplings with runaway potential as the behavior of the potential is the same for all couplings before the end of inflation.
@@ 2.4 Hyperkination and Kination
$3H^{2}=\kappa\rho_{\chi}$ (2.26)
where $H$ is the Hubble parameter. Following the methodology given by [95], in order to eliminate $H$ from the equation of motion, we change the time variable to the number of $e$-folds, $N=\ln{a}$, where $\mathrm{d}N=H\mathrm{d}\tau$. We normalize the scale factor such that, for $a=e^{N}=1$ for $N=0$, $N$ denotes the number of $e$-folds after the end of inflation. Let the overhead bar denote the derivative with respect to $N$ (the prime on $h$ indicates derivative with respect to $\chi$) and let us assume $\dot{\chi}>0$. From (2.26),
@@
We compare the convergence of the derivative of the field with respect to the number of $e$-folds $N$ to the kinetic attractor value for (2.42) and for different exponents of coupling $t$ from (2.32), considering all the constraints on initial conditions and length of hyperkination and transition in Figure 2. The three black dashed vertical lines denote the start of kination, $N_{\text{kin}}$, the end of inflation, the start of hyperkination, $N_{\text{inf}}^{{\text{end}}}=0,$ and the limiting case for length of hyperkination, $N_{\text{hyp}}^{{\text{end}}}=0.2$. The constraint on transition length between hyperkination and kination from (2.48) suggests that the kinetic terms start to dominate slightly before the inflation completely ends and hyperkination lasts for a small duration. The end of hyperkination is followed by the phase of standard kination, where the field rolls forward freely. The epoch of hyperkination is fastest for the initial approximation in (2.42). The length increases as we increase the coupling exponent in (2.32), and convergence is achieved at values mentioned in Table 1. The field derivative for $t=3$ never approaches the attractor value.
@@ 3 Reheating Mechanism
where $b$ and $n$ are small and positive coupling constants. For simplicity, we assume that the bare masses of the fields $\zeta$ and $\psi$ are very small, such that $m_{\zeta}(\chi)=b\chi$ and $m_{\psi}(\chi)=|n\chi|$. We can consider the effective potential of the scalar field after kination as,
@@
Typically, coupling constants of the interaction of the inflaton field with matter are extremely small [110], we consider $b\sim\mathcal{O}(0.1)$ and $n\sim\mathcal{O}(0.01)$. $\lambda$ is the coupling constant for the Higgs-like potential. Considering it is stable, we choose $\lambda\sim\mathcal{O}(0.01)$. The only free parameter now is the mass of the scalar field.
@@
To get an era of reheating, the mass of the inflaton field should be fine-tuned such that the decay rate causes the inflaton field to decay faster than radiation. This happens for $m_{\chi}=10^{4}-10^{5}$ GeV, see the left panel of Figure 3. If no decay process is considered (black dashed line), the scalar field energy density will never be sub-dominant than radiation. As we increase the mass parameter, the energy density dilutes faster, eventually letting radiation take over for $m_{\chi}>8\times 10^{3}$ GeV. Considering that the decay processes take time and do not happen as quickly as for $m_{\chi}=10^{5}$ GeV, we choose the mass of the scalar field to be $10^{4}$ GeV. The length of hyperkination is fixed to the upper bound of $\Delta N_{\text{hyp}}=0.2$; for a choice of $m_{\chi}=10^{4}$ GeV, the length of the kination phase is $\Delta N_{\text{kin}}=4$. The panel on the right of Figure 3 is the new evolution of energy density accounting for the decay process in our alternate history theory for couplings with exponents $t>0$. The reheating temperature can be estimated by $T_{\text{reh}}\sim 0.1\sqrt{\Gamma M_{\text{Pl}}}\sim 10^{8}$ GeV for $m_{\chi}=10^{4}$ GeV giving a successful and efficient mechanism in such scenarios.
@@ 4.1 Background Solutions
where $\epsilon$ is the first Hubble slow-roll parameter. Note that we will denote the Hubble parameter at the end of inflation $H_{\text{end}}$ as $H$. Solving for all the epochs, see Appendix C for detailed derivation, we get
@@ 4.2 Gravitational Wave Mode Functions
where prime denotes derivative with respect to the conformal time $\eta$. From the previous section, we can write,
@@
where $H_{\nu}^{(1)}$ and $H_{\nu}^{(2)}$ are Hankel functions of the first and second kind, respectively. We solve for the mode functions in detail in Appendix D and note the final results here,
@@
To determine the constants $\alpha_{\pm}$, $\beta_{\pm}$, and $\gamma_{\pm}$; we demand the continuity of $v_{k}^{s}$ and its derivative at the transition times $\eta_{\text{end}}$, $\eta_{\text{kin}}$, $\eta_{\text{reh}}$. The details for this mode-matching process are mentioned in Appendix E. The results are as follows,
@@
We check the time-evolution of the perturbations by fixing the length of the period of hyperkination and varying the order of coupling; see Figure 5. We keep $\Delta N_{\text{hyp}}=6.18$ so that we can gauge the changes if there are any from the variation. The order of the coupling decreases as we go from top to bottom. Similar to the results obtained in Figure 4 for kination, we see that the perturbations get a boost in energy for all orders. However, as the order decreases, the amplitude of produced gravitational wave modes during hyperkination increases, indicating that the hyperkination phase will contribute significantly to the primordial background for such couplings. Another feature is the increasing length of the period of kination with decreasing order. This suggests that the energy from GWs during kination will also increase for every decreasing order. In Figure 6, we vary $\Delta N_{\text{hyp}}$ for the $t=2$ coupling. We do not observe any change in GW perturbations generated during hyperkination. However, the kination phase reduces in length, and the modes decay faster with decreasing length of hyperkination.